If c = 18a + 24b, where a and b are positive integers, then c must be divisible by which of the following?
a) 4
b) 6
c) 9
d) 12
e) 72
Than you!!! :-)

Answer by Edwin McCravy(20055)

 18a + 24b

by the distributive principle (used in reverse)

6(3a + 8b)

Therefore it must be divisible by 6

An easy counter-example for the others is when a = 1, and b = 1

c = 18a + 24b = 18(1) + 24(1) = 18 + 24 = 42 which is 

not divisible by 4, 9, 12, or 72

Edwin